OpenStudy (sew01246):

Find equation of the tangent line to the curve x^2+ siny =xy^2 + 1 at (1,0)

OpenStudy (phi):

can you find dy/dx ? (using implicit differentiation) or do you need help ?

OpenStudy (sew01246):

i need help

OpenStudy (phi):

x^2+ siny =xy^2 + 1 I assume you can find the derivative with respect to x of x^2 ?

OpenStudy (phi):

what is \[ \frac{d}{dx} x^2 \]?

OpenStudy (sew01246):

0?

OpenStudy (phi):

the derivative of a constant is 0 use the power rule

OpenStudy (phi):

the power rule is \[ \frac{d}{dx} x^n = n x^{n-1} \frac{dx}{dx} = n x^{n-1}\]

OpenStudy (sew01246):

so how would i put power rule to use?

OpenStudy (phi):

The very first derivatives you learn is the power rule. The problem you posted is much more advanced, and requires implicit differentiation, product rule, and trig functions.

OpenStudy (photon336):

Well, technically you also have to use product rule for differentiation too. I agree with @phi for this problem there's alot going on here. if I were you I wouldn't attempt this until you understand the power rule first. first let's move the other term to the left hand side. \[x^{2}+siny-xy^{2} = 1 \] then differentiate both sides \[\frac{ d }{ dx }(x+\sin(y)-xy^{2}) = \frac{ d }{ dx }(1)\]

OpenStudy (phi):

If you need more background or review, you can try Khan's videos https://www.khanacademy.org/math/calculus-home/differential-calculus/taking-derivatives

OpenStudy (photon336):

But start off with the power rule, and the basic rules for differentiation first.

OpenStudy (photon336):

otherwise it's not going to make sense.

OpenStudy (sew01246):

okay, after I then differentiate both sides

OpenStudy (sew01246):

what's next?

OpenStudy (phi):

Here is a graph of the problem

OpenStudy (sew01246):

Thanks for the video, as you can see I'm struggling with this problem.

OpenStudy (photon336):

what is \[\frac{ d }{ dx } (x^{2}) = ? \]

OpenStudy (photon336):

@sew01246

OpenStudy (sew01246):

0?

OpenStudy (photon336):

Here is the rule \[\frac{ d }{ dx } x^{n} = nx^{n-1}\]

OpenStudy (photon336):

in \[x^{2} \rightarrow what~is~n?\]

OpenStudy (sew01246):

1?

OpenStudy (photon336):

\[\frac{ d }{ dx } x^2 = 2x^{2-1} = 2x\]

OpenStudy (photon336):

@sew01246 can you see why?

OpenStudy (sew01246):

no, can you explain?

OpenStudy (photon336):

x^2 what is the exponent ?